# Do Code LLMs Do Static Analysis?

### v1

@@ Abstract
This paper investigates code LLMs’ capability of static analysis during code intelligence tasks such as code summarization and generation. Code LLMs are now household names for their abilities to do some programming tasks that have heretofore required people. The process that people follow to do programming tasks has long been understood to require static analysis. For example, human programmers navigate the call graph of large programs to comprehend the different parts of those programs. Education in programming includes static analysis under the assumption that better static analysis skills beget better programming. Yet while popular culture is replete with anthropomorphic references such as LLM “reasoning”, in fact code LLMs could exhibit a wholly alien thought process to humans. This paper studies the specific question of static analysis by code LLMs. We use three different static analysis tasks (callgraph generation, AST generation, and dataflow generation) and three different code intelligence tasks (code generation, summarization, and translation) with two different open-source models (Gemini and GPT-4o) and closed-source models (CodeLlaMA and Jam) as our experiments. We found that LLMs show poor performance on static analysis tasks and that pretraining on the static analysis tasks does not generalize to better performance on the code intelligence tasks.
Keywords:
static analysis code intelligence tasks large language modelsLLM reasoning
@@ 1 Introduction
Popular discussion around large language models (LLMs) is replete with anthropomorphic references such as LLM “reasoning” and “creativity” (Tian et al., 2024; Zhao et al., 2024). Even dialogue in scientific literature uses biological terms such as “attention” and “hallucination” to refer to parts of LLM architecture or behavior. The implication is that LLMs think and act like people, with human intentions and human thought processes. One may conclude that this implication is clearly hyperbole – many authors argue that impressive LLM abilities are in fact merely pattern matching across big data input (Wang et al., 2023) and/or carefully curated finetuning (Chen and Mueller, 2024). Yet the processes that LLMs follow are controversial and there is not yet sufficient empirical evidence on specific LLM behaviors to draw strong conclusions.
@@
One group of LLM behaviors in the spotlight is in software development tasks. Code LLMs are language models specializing in these tasks such as generating code samples from natural language prompts, code comment generation, and translation from one programming language to another. Many programmers use code LLMs every day as a normal part of their workflow. For example, Google estimated in November of 2024 that over 25% of its new code is written by AI and later reviewed by engineers (arstechnica, 2024). The rapid rise of LLM-generated code has led to anxiety about the safety and security of programs containing that code (Mohsin et al., 2024; Wang et al., 2024a), privacy and other legal issues (He et al., 2024), and even the future of software engineering as a career (Rasnayaka et al., 2024; Tona et al., 2024). Knowing how code LLMs do coding tasks could shed light on these concerns.
@@
Meanwhile, a fundamental part of what people do during programming tasks is static analysis. Static analysis is the process of examining code without executing it (Ayewah et al., 2008). Studies of human programmers have for decades shown how people build a mental model of program behavior through static tracing of three relationships in particular: function invocations (i.e., function call relationships), data flow among variables, and the hierarchical structure of code elements, such as how expressions, statements, and blocks are organized (Littman et al., 1987). A rich body of literature describes these and similar static analysis tasks which underpin development of static analysis tools built into popular IDEs (Luo et al., 2021).
@@
This paper asks the question: do code LLMs do static analysis? Humans do static analysis as a core part of software development, so if code LLMs follow a process similar to humans, then one would expect code LLMs to also do static analysis. We conduct experiments in which we 1) observe the performance of LLMs on software development tasks, 2) observe the performance of those same LLMs on static analysis tasks, 3) finetune open-source LLMs to do static analysis tasks, and 4) observe if the finetuned LLMs’ performance on development tasks changes. One would expect improved performance on development tasks if the code LLMs are trained to do static analysis better.
@@
We found that code LLMs generally do a poor job at static analysis tasks but have strong performance at coding tasks. We also found that finetuning LLMs with static analysis tasks improved models’ ability to do simple static analysis tasks, but did not improve models’ ability on more complicated static analysis and coding tasks. We found these results over two open-source (CodeLlaMA, jam) and two closed-source models (GPT-4o, Gemini) and two programming languages (Java, C/C++). We studied three static analysis tasks (call graph, data flow graph, and abstract syntax tree generation) and three programming tasks (code summarization, code translation, and code generation). Our findings contribute to the ongoing academic debate about how LLMs function and have implications for task generalization, training procedures, and the design of future LLMs for software engineering tasks.
@@ 2.1 Static Analysis in Program Comprehension
Empirical Studies Program comprehension has been a target of empirical studies for a decade. Several empirical studies have been conducted to understand how programmers comprehend code. Littman et al. (1987) found that programmers use data flow for comprehending code. Pennington (1987) also found that programmers combine data flow and control flow to comprehend source code in their study with 80 professional programmers. Shneiderman and Mayer (1979) have similar conclusion that syntactic information is as important as semantic information. The way that programmers represent the code mentally can be referred as mental models (Siegmund, 2016; O’brien, 2003; Storey, 2005; Harth and Dugerdil, 2017)Harth and Dugerdil (2017) classified this into three different periods: 1) classical period 2) optimistic period 3) pragmatic period. In the classical period, several different strategies on how programmers understand code were proposed. During the optimistic period, several tools and studies were developed and conducted. The pragmatic period focus more on measurement instead of creating new methods.
@@
Static Analysis Various studies have established that three static analysis tasks in particular are crucial to program comprehension: 1) call graph analysis, 2) data flow analysis, and 3) Abstract Syntax Tree (AST) generation. Jiang et al. (2017) showed that programmers use call graphs to do change impact analysis. Konopka et al. (2018) conducted an empirical study with novice and intermediate programmers and showed the importance of using data flow analysis for program comprehension tasks. Another study conducted by  Maalej et al. (2014) also showed that some programmers drew the data flow and control flow chart during program comprehension. Hovemeyer et al. (2016) used ASTs to analyze students’ submission to help students on their programming tasks. Overall, these show that programmers rely on static analysis to do program comprehension tasks.
@@
Static analysis helps programmers analyze code base without the need to execute the source code and any input. Different theories on static analysis have been proposed. Park et al. (2021) proposed the parametric static analysis to theorize the static analysis. Static analysis helps programmers in different aspects of software developments (Vassallo et al., 2020; Beller et al., 2016)Vassallo et al. (2018) found that developers use the warning message from automatic static analysis tools to find the bugs in the source code. Ayewah et al. (2008) combined the static analysis tools for bug detection. Static analysis can be used to aid the program comprehension (Eisenbarth et al., 2001). In the study conducted by Tymchuk et al. (2018), they showed that static analysis tools can also be served as an educational tool. Singh et al. (2017) showed that static analysis tools can reduce the workload of code review.
@@ 2.2 Code Large Language Models
Code LLMs have been the trend of in the recent year (Zhang et al., 2023; Jiang et al., 2024a; Chen et al., 2025) and have been applied to several different software engineering tasks (Nam et al., 2024; Gu, 2023; Zheng et al., 2025; He et al., 2025; Zhong and Wang, 2024). In this section, we discuss the history of code LLMs. Specifically, we focus on how LLMs are trained and how the source code is represented during different stages. Then, we discuss the application of LLMs in software engineering and studies on the capabilities of code LLMs.
@@
History Sun et al. (2024a) divided the history of code LLMs into three groups: 1) neural language models for code 2) code pre-trained models 3) code LLMs. Natural language models represent the code as the natural language. Different methods have been proposed to represent code. Alon et al. (2019) encode each code token as a word vector. Zhang et al. (2019) showed that ASTs can better encode the information in the code. LeClair et al. (2020) applied graph neural network to represent code for code summarization. Code pre-trained models pretrain the language models with huge amount of source code and finetuned it with downstream tasks. Wang et al. (2021) pretrained the models with encoder-decoder transformer architectures. Li et al. (2023b) proposed a decoder architecture only model for pretrainig. More recently, the attention shift toward scaling up the models and in-context learning. The models include open source models Codellama (Roziere et al., 2023) and several closed source models (Brown et al., 2020; Anil et al., 2023).
@@
Applications LLMs have been applied to different domains of software engineering tasks (Zheng et al., 2025, 2023; Jin et al., 2024)Sarsa et al. (2022) and Sun et al. (2023b) examine the capabilities of LLMs on code summarization. Khan and Uddin (2022) applied LLMs to generate the summarization. Koziolek et al. (2024) leveraged the LLMs and retrieval augmentation for code generations. Different evaluation methods have also been studied (Liu et al., 2024c, b)Yuan et al. (2024) proposed multi-agent systems for code translation. Pan et al. (2023) applied decoder-only LLMs to translate different programming languages into Python. In addition to program comprehension tasks, Xue et al. (2024) used LLMs to generate the test case for FinTech software testing. Lu et al. (2024) use LLMs for vulnerabilities detection with in-context learning. Yin et al. (2024) and a LLM-based framework for automatic program repair. Overall, LLMs are widely used in the software engineering.
@@
Capability Studies Several studies on LLMs capability have been published (Yang et al., 2024; Li et al., 2024b; Jiang et al., 2024a; Fan et al., 2024)Yang et al. (2024) divided the capabilities of code LLMs into three different ares 1) programming skills 2) complex reasoning 3) structure knowledge. Li et al. (2022) showed that LLMs defeat around half of the programmers during programming competitions. Several studies have showed the promising results on using multiple LLMs as an agent for code generation (Liu et al., 2023a, b; Talebirad and Nadiri, 2023)Li et al. (2023a) showed that LLMs can do complex reasoning on debugging. Chen et al. (2022) showed that LLMs can reason the output of the code for several math problems. Although several papers showed promising results on reasoning, several studies have also showed that LLMs can only do reasoning when the code is simple (Liu et al., 2024a; Chen et al., 2024)Madaan et al. (2022) and Wang et al. (2022) showed the ability of LLMs for structure prediction after few-shot learning.
@@ 3 Experimental Design
Our experiments center around three topics: 1) we observe LLM performance on code development tasks, 2) we observe LLM performance on static analysis tasks in baseline, in-context learning setting, and finetuned setting, and 3) we finetune LLMs to do static analysis and determine if those models’ performance on code development tasks increases. This section covers our research questions, methods to answer those questions, and experimental settings.
@@ 3.1 Research Questions
Our research objective is to determine if LLMs that do code development tasks also “know how” to do static analysis and to determine if models’ ability to do static analysis is connected to the models’ ability to do code development tasks. We ask the following Research Questions (RQs):
@@
What is the performance of open-source LLMs on code development tasks when finetuned for those tasks under recommended conditions?
@@
The rationale for RQ1 is to establish a baseline in a controlled environment of how well representative LLMs perform on code development tasks with common datasets and settings. Using open-source LLMs allows us to avoid data contamination issues that are present with these datasets in commercial LLMs (Jiang et al., 2024b). The rationale behind RQ2 is that state-of-the-art commercial LLMs are already known to perform code development tasks well (Zheng et al., 2025), but we do not yet know how well these models perform static analysis tasks. The rationale behind RQ3 is that LLMs may be able to perform static analysis tasks if they are explicitly trained to do so, and the availability of open source models provides us the opportunity to do this training. Finally, the rationale behind RQ4 is that models may do better at code development tasks if they can also do static analysis tasks because static analysis tasks are often a part of code development tasks.
@@ 3.2 Code Development Tasks
Code Summarization Code summarization is the task of writing natural language descriptions of a section of code (Haiduc et al., 2010). Code summarization is important because it is the core of automatic documentation generation and related tool support for helping programmers understand code; the history of research efforts in code summarization has been chronicled in recent surveys (Zhang et al., 2024b, 2022). We study this task in this paper because code summarization has become an advertised feature of code LLMs (Sun et al., 2024b) and because different studies show how people do static analysis in order to do code summarization (Rodeghero and McMillan, 2015; Stapleton et al., 2020). In this paper, for Java we use the “170k” dataset by Su and McMillan (2024). That dataset has its origins in a study of code summarization dataset recommendations by LeClair and McMillan (2019) and further vetted by (Bansal et al., 2021). A major advantage to the dataset is that it’s test set corresponds to a set held out of training in the jam model, which strongly avoids data contamination concerns (see Section 3.4). For C/C++ we use a dataset by Liu et al. (2021) which followed similar best practices (e.g., (LeClair and McMillan, 2019; Allamanis, 2019)).
@@
Code Translation Code translation is the task of writing a program in one language that has the same functionality as a program in another programming language. Code translation can help program comprehension by in converting pseudo code into machine readable code (Holt and Turanski, 1960) and translate less-readable programming language into a more human-readable programming language (Hong and Ryu, 2024). Several studies have shown that code translation requires static analysis when done manually by people (Siddiqui et al., 2023; Feautrier, 1991). In this paper, we use code translation dataset in CodeXGLUE (Lu et al., 2021) as our dataset for Java. CodeXGLUE collects several datasets for code intelligence tasks. CodeXGLUE divides the datasets into code-code, code-text, text-code, and text-code. The code translation dataset that we use is collected from several public repositories with deduplication and clear train/test/validation separation. For C/C++, we use the C to Rust dataset in Huggingface website (Yu, 2023).
@@
Code Generation Code generation is the task of generating source code that implements the functionality described in natural language. Code generation can be used as an educational tool to help novice programmers learn programming (Becker et al., 2023; Mosterman, 2006) and is widely used to improve the productivity of software developers (Ross et al., 2023; Xu et al., 2022). Static analysis is part of generating code in that an understanding of the static structure leads to higher code quality and few bugs (Barke et al., 2023; Nachtigall et al., 2019). In this paper, we used CONCODE datset (Iyer et al., 2018) from CodeXGLUE Lu et al. (2021) for Java. The dataset is collected from 33,000 repositories and they filtered out the methods without Javadoc as natural language description. Although the dataset has very clear train/validation/test split, we use separate half of the validation set as our testset because we do not find the ground-truth for testset at the time we download. For C/C++, we use Xlcost (Zhu et al., 2022). Xlcost is a dataset that contains seven different programming languages and natural language, though we only use the C/C++ component.
@@
Development Code Summarization Java 170k  Su and McMillan (2024) x x
C/C++ Liu et al. (2021) Liu et al. (2021) x
Code Translation Java CodeXGLUE Lu et al. (2021) x x
C/C++ C to Rust  Yu (2023) x
Code Generation Java CONCODE Lu et al. (2021) x x
C/C++ Xlcost  Zhu et al. (2022) x
Static Analysis AST Generation Java funcom  LeClair and McMillan (2019) x x
C/C++  Haque et al. (2021)  Haque et al. (2021) x
Call Graph Generation Java funcom  LeClair and McMillan (2019) x x
C/C++  Haque et al. (2021)  Haque et al. (2021) x
Dataflow Graph Generation Java funcom  LeClair and McMillan (2019) x x
C/C++  Haque et al. (2021)  Haque et al. (2021) x

### v2

@@ Abstract
This paper investigates code LLMs’ capability of static analysis during code intelligence tasks such as code summarization and generation. Code LLMs are now household names for their abilities to do some programming tasks that have heretofore required people. The process that people follow to do programming tasks has long been understood to require static analysis. For example, human programmers navigate the call graph of large programs to comprehend the different parts of those programs. Education in programming includes static analysis under the assumption that better static analysis skills beget better programming. While popular culture is replete with anthropomorphic references such as LLM “reasoning”, in fact code LLMs could exhibit a wholly alien thought process to humans. This paper studies the specific question of static analysis by code LLMs. We use three different static analysis tasks (callgraph generation, AST generation, and dataflow generation) and three different code intelligence tasks (code generation, summarization, and translation) with two different open-source models (Gemini and GPT-4o) and closed-source models (CodeLlaMA and Jam) as our experiments. We found that LLMs show poor performance on static analysis tasks and that pretraining on the static analysis tasks does not generalize to better performance on the code intelligence tasks and vice versa.
@@ 1 Introduction
Popular discussion around large language models (LLMs) is replete with anthropomorphic references such as LLM “reasoning” and “creativity” (tian2024thinking; zhao2024assessing). Even dialogue in scientific literature uses biological terms such as “attention” and “hallucination” to refer to parts of LLM architecture or behavior. The implication is that LLMs think and act like people, with human intentions and human thought processes. One may conclude that this implication is clearly hyperbole – many authors argue that impressive LLM abilities are in fact merely pattern matching across big data input (wang2023can) and/or carefully curated finetuning (chen2024automated). Yet the processes that LLMs follow are controversial and there is not yet sufficient empirical evidence on specific LLM behaviors to draw strong conclusions.
@@
One group of LLM behaviors in the spotlight is in software development tasks. Code LLMs are language models specializing in these tasks such as generating code samples from natural language prompts, code comment generation, and translation from one programming language to another. Many programmers use code LLMs every day as a normal part of their workflow. For example, Google estimated in November of 2024 that over 25% of its new code is written by AI and later reviewed by engineers (arstechnica2024google). The rapid rise of LLM-generated code has led to anxiety about the safety and security of programs containing that code (mohsin2024can; wang2024your), privacy and other legal issues (he2024emerged), and even the future of software engineering as a career (rasnayaka2024empirical; tona2024exploring). Knowing how code LLMs do coding tasks could shed light on these concerns.
@@
Meanwhile, a fundamental part of what people do during programming tasks is static analysis. Static analysis is the process of examining code without executing it (ayewah2008using). Studies of human programmers have for decades shown how people build a mental model of program behavior through static tracing of three relationships in particular: function invocations (i.e., function call relationships), data flow among variables, and the hierarchical structure of code elements, such as how expressions, statements, and blocks are organized (littman1987mental). A rich body of literature describes these and similar static analysis tasks which underpin development of static analysis tools built into popular IDEs (luo2021ide).
@@
This paper asks the question: do code LLMs do static analysis? Humans do static analysis as a core part of software development, so if code LLMs follow a process similar to humans, then one would expect code LLMs to do static analysis. We conduct experiments in which we 1) observe the performance of LLMs on software development tasks, 2) observe the performance of those same LLMs on static analysis tasks, 3) finetune open-source LLMs to do static analysis tasks, 4) observe if the finetuned LLMs’ performance on development tasks changes, and 5) observe if LLMs have the byproduct as the results of being trained on code intelligence tasks. One would expect improved performance on development tasks if the code LLMs are trained to do static analysis better and have static analysis as a byproduct.
@@
We found that code LLMs generally do a poor job at static analysis tasks but have strong performance at coding tasks. We also found that finetuning LLMs with static analysis tasks improved models’ ability to do simple static analysis tasks, but did not improve models’ ability on more complicated static analysis and coding tasks. In addition, training LLMs on code intelligence tasks does not lead to better generation of static analysis tasks. We found these results over two open-source (CodeLlaMA, jam) and two closed-source models (GPT-4o, Gemini) and two programming languages (Java, C/C++). We studied three static analysis tasks (call graph, data flow graph, and abstract syntax tree generation) and three programming tasks (code summarization, code translation, and code generation). Our findings contribute to the ongoing academic debate about how LLMs function and have implications for task generalization, training procedures, and the design of future LLMs for software engineering tasks.
@@ 2.1 Static Analysis in Program Comprehension
Empirical Studies Program comprehension has been a target of empirical studies for a decade. Several empirical studies have been conducted to understand how programmers comprehend code. littman1987mental found that programmers use data flow for comprehending code. pennington1987stimulus also found that programmers combine data flow and control flow to comprehend source code in their study with 80 professional programmers. shneiderman1979syntactic have similar conclusion that syntactic information is as important as semantic information. The way that programmers represent the code mentally can be referred as mental models (siegmund2016program; o2003software; storey2005theories; harth2017program)harth2017program classified this into three different periods: 1) classical period 2) optimistic period 3) pragmatic period. In the classical period, several different strategies on how programmers understand code were proposed. During the optimistic period, several tools and studies were developed and conducted. The pragmatic period focus more on measurement instead of creating new methods.
@@
Static Analysis Various studies have established that three static analysis tasks in particular are crucial to program comprehension: 1) call graph analysis, 2) data flow analysis, and 3) Abstract Syntax Tree (AST) generation. For example, jiang2017programmers conducted extensive studies on impact analysis tasks. They recruited nine professional programmers for a depth study and 35 programmers for the breath study. They found that callgraphs help programmers to navigate the source code. wallace2025programmer conducted a project-level eye-tracking study with 10 Java programmers for source code summarization. They suggested using callgraph to develop context-aware LLMs for source code summarization. konopka2018data conducted an empirical study with novice and intermediate programmers and showed the importance of using data flow analysis for program comprehension tasks. In addition, they proposed a data-flow based metrics to evaluate how human programmers comprehend a program. bergeretti1985information also show that the information flow helps various stages of software development. maalej2014comprehension conducted a study on how software developers comprehend source code with 28 software developers. They found that some participants drew the data flow to comprehend the program. hovemeyer2016control showed that simple AST helps to understand the strategies that students use for their programming assignments. agrahari2020astar show that AST helps the novice programmers to learn the fundamental of data structure. schanzer2019accessible developed an AST-based tool for vision impaired programmers. They found the improvement with their tools. Overall, these show that programmers rely on static analysis to do program comprehension tasks.
@@
Static analysis helps programmers analyze code base without the need to execute the source code and any input. Different theories on static analysis have been proposed. park2021survey proposed the parametric static analysis to theorize the static analysis. Static analysis helps programmers in different aspects of software development (vassallo2020developers; beller2016analyzing)vassallo2018context found that developers use the warning message from automatic static analysis tools to find the bugs in the source code. ayewah2008using combined the static analysis tools for bug detection. Static analysis can be used to aid the program comprehension (eisenbarth2001aiding). In the study conducted by tymchuk2018feedback, they showed that static analysis tools can also be served as an educational tool. singh2017evaluating showed that static analysis tools can reduce the workload of code review.
@@ 2.2 Code Large Language Models
Code LLMs have been the trend of in the recent year (zhang2023survey; jiang2024survey; chen2025an) and have been applied to several different software engineering tasks (nam2024using; gu2023llm; zheng2025towards; he2025llmbased; zhong2024can). In this section, we discuss the history of code LLMs. Specifically, we focus on how LLMs are trained and how the source code is represented during different stages. Then, we discuss the application of LLMs in software engineering and studies on the capabilities of code LLMs.
@@
History sun2024survey divided the history of code LLMs into three groups: 1) neural language models for code 2) code pre-trained models 3) code LLMs. Natural language models represent the code as the natural language. Different methods have been proposed to represent code. alon2019code2vec encode each code token as a word vector. Zhang2019Novel showed that ASTs can better encode the information in the code. leclair2020improved applied graph neural network to represent code for code summarization. Code pre-trained models pretrain the language models with huge amount of source code and finetuned it with downstream tasks. wang2021codet5 pretrained the models with encoder-decoder transformer architectures. li2023starcoder proposed a decoder architecture only model for pretrainig. More recently, the attention shift toward scaling up the models and in-context learning. The models include open source models Codellama (roziere2023code) and several closed source models (brown2020language; anil2023palm).
@@
Applications LLMs have been applied to different domains of software engineering tasks (zheng2025towards; zheng2023survey; jin2024llms)sarsa2022automatic and sun2023automatic examine the capabilities of LLMs on code summarization. khan2022automatic applied LLMs to generate the summarization. Koziolek2024llmbased leveraged the LLMs and retrieval augmentation for code generations. Different evaluation methods have also been studied (liu2024your; liu2024exploring)yuan2024transagent proposed multi-agent systems for code translation. pan2023stelocoder applied decoder-only LLMs to translate different programming languages into Python. In addition to program comprehension tasks, Xue2024llm4fin used LLMs to generate the test case for FinTech software testing. lu2024grace use LLMs for vulnerabilities detection with in-context learning. yin2024thinkrepair and a LLM-based framework for automatic program repair. Overall, LLMs are widely used in the software engineering.
@@
Capability Studies Several studies on LLMs capability have been published (yang2024if; li2024personal; jiang2024survey; fan2024exploring)yang2024if divided the capabilities of code LLMs into three different ares 1) programming skills 2) complex reasoning 3) structure knowledge. li2022competition showed that LLMs defeat around half of the programmers during programming competitions. Several studies have showed the promising results on using multiple LLMs as an agent for code generation (liu2023dynamic; liu2310dynamic; talebirad2023multi)li2023hitchhiker showed that LLMs can do complex reasoning on debugging. chen2022program showed that LLMs can reason the output of the code for several math problems. Although several papers showed promising results on reasoning, several studies have also showed that LLMs can only do reasoning when the code is simple (liu2024codemind; chen2024evaluating)madaan2022language and wang2022code4struct showed the ability of LLMs for structure prediction after few-shot learning.
@@ 3 Experimental Design
Our experiments center around four topics: 1) we observe LLM performance on code development tasks, 2) we observe LLM performance on static analysis tasks in baseline, in-context learning setting, and finetuned setting, 3) we finetune LLMs to do static analysis and determine if those models’ performance on code development tasks increases, and 4) we observe whether LLMs can do static analysis as a byproduct of being trained on code development tasks. This section covers our research questions, methods to answer those questions, and experimental settings.
@@ 3.1 Research Questions
Our research objective is to determine if LLMs that do code development tasks also “know how” to do static analysis and to determine if models’ ability to do static analysis is connected to the models’ ability to do code development tasks and vice versa. We ask the following Research Questions (RQs):
@@
What is the performance of open-source LLMs on static analysis tasks after they have been finetuned for code intelligence tasks?
  RQ5
What are the common error types in static analysis?
The rationale behind RQ1 is that state-of-the-art commercial LLMs are already known to perform code development tasks well (zheng2025towards), but we do not yet know how well these models perform static analysis tasks. The rationale behind RQ2 is that LLMs may be able to perform static analysis tasks if they are explicitly trained to do so, and the availability of open-source models provides us the opportunity to do this training. The rationale behind RQ3 is to examine whether finetuning the models with static analysis tasks could improve the performance of code intelligence tasks. This is because human programmers use various static analysis tasks as an intermediate thought process for code intelligence tasks (Refer to Section 2.1). If a model truly has the anthropomorphic characteristic, static analysis should provide the fundamental knowledge for LLMs to perform code intelligence tasks. The rationale for RQ4 is that models can internalize some aspects of static analysis after finetuning models on code intelligence tasks as a byproduct if models can do static analysis tasks. This research question allows us to have a more thorough evaluation on whether LLMs can do static analysis tasks. The rationale for RQ5 is that LLMs may fall short on specific areas. RQ5 provides the insights on where the LLMs fall short on static analysis tasks.
@@ 3.2 Code Development Tasks
Code Summarization Code summarization is the task of writing natural language descriptions of a section of code (haiduc2010on). Code summarization is important because it is the core of automatic documentation generation and related tool support for helping programmers understand code; the history of research efforts in code summarization has been chronicled in recent surveys (zhang2024review; zhang2022survey). We study this task in this paper because code summarization has become an advertised feature of code LLMs (sun2024source) and because different studies show how people do static analysis in order to do code summarization (rodeghero2015empirical; stapleton2020human). In this paper, for Java we use the “170k” dataset by su2024distilled. That dataset has its origins in a study of code summarization dataset recommendations by leclair2019recommendations and further vetted by (bansal2021project). A major advantage to the dataset is that it’s test set corresponds to a set held out of training in the jam model, which strongly avoids data contamination concerns (see Section 3.4). For C/C++ we use a dataset by liu2021retrievalaugmented which followed similar best practices (e.g., (leclair2019recommendations; allamanis2019adverse)).
@@
Code Translation Code translation is the task of writing a program in one language that has the same functionality as a program in another programming language. Code translation can help program comprehension by in converting pseudo code into machine readable code (holt1960man) and translate less-readable programming language into a more human-readable programming language (hong2024dont). Several studies have shown that code translation requires static analysis when done manually by people (siddiqui2023towards; feautrier1991dataflow). In this paper, we use code translation dataset in CodeXGLUE (lu2021codexglue) as our dataset for Java. CodeXGLUE collects several datasets for code intelligence tasks. CodeXGLUE divides the datasets into code-code, code-text, text-code, and text-code. The code translation dataset that we use is collected from several public repositories with deduplication and clear train/test/validation separation. For C/C++, we use the C to Rust dataset in Huggingface website (ccodetrandataset). For execution-based evaluation, we finetuned our models with CodeTransOcean (yan2023codetransocean) and test our models with datasets proposed by pan2024lost because we do not find test cases in CodeTransOcean at the time we download the dataset.
@@
Code Generation Code generation is the task of generating source code that implements the functionality described in natural language. Code generation can be used as an educational tool to help novice programmers learn programming (becker2023programming; mosterman2006automatic) and is widely used to improve the productivity of software developers (ross2023programmer; xu2022inide). Static analysis is part of generating code in that an understanding of the static structure leads to higher code quality and few bugs (barke2023grounded; nachtigall2019explaining). In this paper, we used CONCODE dataset (iyer2018mapping) from CodeXGLUE (lu2021codexglue) for Java. The dataset is collected from 33,000 repositories and they filtered out the methods without Javadoc as natural language description. Although the dataset has very clear train/validation/test split, we use separate half of the validation set as our testset because we do not find the ground-truth for testset at the time we download. For C/C++, we use Xlcost (zhu2022xlcost). Xlcost is a dataset that contains seven different programming languages and natural language, though we only use the C/C++ component. For execution-based evaluation, we finetuned our models with the programming competition dataset proposed by li2022competition and evaluated the results with pass@k and compilation rate. Note that the dataset includes the problems with several programming languages. We only used problems that have C/C++ solutions. We explain the metrics that we used in Section 3.5.
@@
Development Code Summarization Java 170k su2024distilled x x
C/C++ liu2021retrievalaugmented liu2021retrievalaugmented x
Code Translation Java CodeXGLUE lu2021codexglue x x
C/C++ C to Rust ccodetrandataset x
Java Codetransocean yan2023codetransocean x
Java PLTranslationEmpirical pan2024lost x
Code Generation Java CONCODE lu2021codexglue x x
C/C++ Xlcost zhu2022xlcost x
C/C++ Code Contest li2022competition x
Static Analysis AST Generation Java funcom leclair2019recommendations x x
C/C++ haque2021action haque2021action x
Call Graph Generation Java funcom leclair2019recommendations x x
C/C++ haque2021action haque2021action x
Dataflow Graph Generation Java funcom leclair2019recommendations x x
C/C++ haque2021action haque2021action x